- Status Report
- Oct 2, 2022
NOAA GOES SATOPS Morning Report 2 March 2001
SATOPS MORNING REPORT – March 2, 2001
No change in the status of the POES spacecraft.
NOAA-14 rev 31793 / F at 1303Z on March: LAC5B was noisy from
spacecraft with 94% of the data recovered. Recorder management prevented
replay from spacecraft.
NOAA-16 AMSUA1 and AMSU-A2 testing will occur on Monday, March 5, through
Thursday, March 8. On March 5, Cold Calibration Position 3 will be
selected for both the AMSU-A1 and AMSU-A2 as the calibration target.
On March 8, Cold Calibration Position 4 will be selected for both instruments
as the calibration target. Questions regarding this test should be
referred to Dong Han at Dong.Han@noaa.gov.
NOAA-16 SBUV testing will occur on Tuesday, March 6 and Wednesday, March
7. On March 6, Flex Memory Segment 2 (TOMS wavelengths) will be selected
at 1215Z. On March 7, at 1213z, Flex Memory Segment 3 (Cloud Cover
Radiometer wavelengths) will be selected. The normal SBUV wavelengths
(flex memory segment 0) will re re-selected at 2023Z on March 7.
Questions regarding this test should be referred to Dong Han at Dong.Han@noaa.gov.
POES operations were nominal over the past 24 hours.
No change in the status of the DMSP spacecraft.
At 1845Z on March 1, a Cat-III data loss was declared on OPS48 for rev
20147 / F (pass time was 0038Z to 0051Z on March 1). AFWA stated
that their data set was short. Data lost: 600 spacecraft seconds/15
playback seconds. Conversations with AFWA continue to determine the
status of this data set.
DMSP operations were nominal over the past 24 hours.
No change in the status of the GOES spacecraft.
All eclipse ops nominal.
To review the eclipse schedules for the GOES-8 and GOES-10 spacecraft,
refer to: http://www.ssd.noaa.gov/SSD/GOES/eclipse.html to identify images
cancelled to satisfy eclipse or instrument keep-out-zone (KOZ) requirements.
GOES operations were nominal over the past 24 hours.
Questions or comments to email@example.com